OpenStudy (anonymous):

Eric purchased three music CDs and a sweatshirt from a band's online store and received a 20% discount. He paid $100 for his purchase. The following week, the same store offered a 40% discount on all its products. Eric's friend Neil took advantage of the sale and bought four music CDs and two sweatshirts for $120 Assuming that the actual prices of the music CDs and the sweatshirts are unchanged, the price of a music CD is $ a. 20 b. 25 c. 30 d. 50 and the price of a sweatshirt is $ a. 20 b. 25 c. 30 d. 50

OpenStudy (anonymous):

anybody?

OpenStudy (anonymous):

It's A for both. Can you fan and medal if correct?

OpenStudy (anonymous):

u sure ?jeje @ciara9271

OpenStudy (anonymous):

one sec ill recheck

OpenStudy (xapproachesinfinity):

eh let x be the price for one cd and y the price for a sweatshirt so Eric paid (3x+y)-0.2(3x+y)=100 the same thing with Neil (4x+2y)-0.4(4x+2y)=120

OpenStudy (xapproachesinfinity):

solve the system

OpenStudy (anonymous):

A&B

OpenStudy (xapproachesinfinity):

well there are other ways to think about this!

OpenStudy (anonymous):

so i subtitute the x and y to se if true?

OpenStudy (xapproachesinfinity):

realize that if the price of the cds or sweatshirt changed when Neil made his purchase this equations would make any sense

OpenStudy (xapproachesinfinity):

How do you solve a system of equations? there are many methods

OpenStudy (anonymous):

working on it jeje ...

OpenStudy (xapproachesinfinity):

Good luck, make sure you do it correctly

OpenStudy (anonymous):

i dont know i give up -_-...i am not really good at this

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

hehe you can't solve a system of two equations? 2.4x+0.8y=100 and 2.4x+1.2y=120

OpenStudy (anonymous):

so it woulbe x=25 and y=50 i guess?

OpenStudy (anonymous):

its hard :P lol

OpenStudy (anonymous):

B & D

OpenStudy (xapproachesinfinity):

just subtract first equation from second you 1.2y-0.8y=20 0.4y=20 ===>y=50 then x=25

OpenStudy (xapproachesinfinity):

you got that right

OpenStudy (xapproachesinfinity):

yes B and D

OpenStudy (anonymous):

is hard the part you have to set up the equation but once u have the equation is easier..but Thakn u so much

OpenStudy (xapproachesinfinity):

You are welcome! eh it is not actually hard, this something you do in your everyday life